The apparent weight of an object increases in an elevator while accelerating upward. A moongphaliwala sells his moongphali using a beam balance in elevator. Will he gain more if the elevator is accelerating up ?

Grade:12

2 Answers

Sparsh gupta
6 years ago
Yes there will be increase in weight..Consider his weight is mass × acceleration due to gravity(m×g)That is acting downwards And now according to newtons third law of motion "every action has an equal and opposite reaction" therefore there will be an reaction in the opposite side And as we knowNet force on a body is equal to mass × acceleration of the body There fore R-(m×g)= m×a R= m×a + m×gR=m(a+g)Hence when a lift accelerates upward there will be increase in weight
Shivakshi
5 years ago
The apparent weight of an object increases in an elevator while accelerating upward. Since the vendor is using a beam balance  therefore the effect will be same on both sides